# references.bib

### v4

@@ Abstract
Artificial intelligence techniques are increasingly being applied to solve control problems, but often rely on black-box methods without transparent output generation. To improve the interpretability and transparency in control systems, models can be defined as white-box symbolic policies described by mathematical expressions. While current approaches to learn symbolic policies focus on static policies that directly map observations to control signals, these may fail in partially observable and volatile environments. We instead consider dynamic symbolic policies with memory, optimised with genetic programming. The resulting policies are robust, and consist of easy to interpret coupled differential equations. Our results show that dynamic symbolic policies compare with black-box policies on a variety of control tasks. Furthermore, the benefit of the memory in dynamic policies is demonstrated on experiments where static policies fall short. Overall, we present a method for evolving high-performing symbolic policies that offer interpretability and transparency, which lacks in black-box models.
@@ 1 Introduction
Many problems in our increasingly complex society can be viewed as control problems, where the goal is to make optimal decisions in real time such as to regulate the behaviour of the controlled system, ranging from smart infrastructure systems to healthcare systems [Annaswamy2023]. In recent years, there have been many successful applications of artificial intelligence (AI) to control problems, offering automated solutions to regulate complex non-linear processes. While the employed control algorithms achieve high performance, they usually require the use of black-box models such as neural networks [adadi2018peeking]. Such black-box models often allow for highly general computations, but lack interpretability of their functioning, making it difficult to understand the computed decisions and actions [loyola2019black]. Transparency is particularly important in control systems, because it aids in identification of biases, fault detection and improvement and could therefore raise confidence of users in the system’s reliability [zhang2021survey].
@@
In contrast to black-box models, white-box models allow for easy inspection of their internal mechanisms [adadi2018peeking]. However, white-box models show an accuracy-interpretability trade-off and there are currently no general training methods to produce white-box approaches that can compete with black-box approaches in terms of efficiency and accuracy [loyola2019black, puiutta2020explainable]. Most white-box approaches opt for post-training model distillation to produce a transparent model that learns the input-output mapping of a black-box model [gou2021knowledge]. Disadvantages of model distillation are that it requires a two-phase learning process and information loss could occur in the distillation phase [ras2022explainable]. Therefore, it may be more desirable to learn white-box models directly. When training an inherently white-box model, interpretability of decisions already occurs at intermediate stages, showing transparent learning, and outputs are directly dependent on explainable features [ras2022explainable].
@@
A promising candidate for white-box control systems are symbolic policies [calegari2020integration, landajuela2021discovering, gou2021knowledge]. Symbolic policies are described by interpretable symbolic expressions, consisting of mathematical operators and state variables. Similar to symbolic regression [bongard2007automated], symbolic policy learning does not require a predefined model structure, therefore it is less affected by human bias or knowledge gaps compared to other data-driven approaches. Additionally, symbolic policies may capture rules that generalise better to other settings, and learned equations tend to be more efficient than neural networks in terms of parameter count [landajuela2021discovering].
Current approaches of symbolic policy learning are limited to learning static policies that have a fixed mapping from observations to control outputs [guo2024efficient, hein2018interpretable, nadizar2024naturally, landajuela2021discovering, gou2021knowledge, shimooka2000generating]. However, extending static policies to real-world scenarios is difficult, because environments can be volatile (change across time) or partially observable (providing noisy and limited information about the state of the environment). Dynamic policies could in principle solve these issues by incorporating memory to integrate past observations [heess2015memory], but it is as of yet unclear how to optimise such policies in closed-loop control tasks, while incorporating interpretability.
@@
Due to the non-differentiable structure of mathematical expressions, gradient-based methods cannot be used for the optimisation of symbolic policies. A viable alternative is genetic programming (GP), an evolutionary algorithm that learn computer programs [koza1994genetic]. GP has seen great successes in symbolic regression problems, which seek to discover interpretable mathematical models describing data [cao2000evolutionary, sakamoto2001inferring, iba2008inference, schmidt2009distilling, quade2016prediction, bongard2007automated, wang2019symbolic]. Therefore we consider GP as a promising approach to optimise dynamic symbolic policies to solve control tasks [hein2018interpretable, nadizar2024naturally].
In the dynamic policies, the GP learns a symbolic representation of a state equation, describing the flow of a differential equation that takes observations as input and generates controls as output. Here, the latent state of the underlying differential equations functions as a memory system to capture long term dependencies in the data, which could improve the policy’s robustness and generalisability [heess2015memory]. This is similar to the use of neural differential equations to estimate parameterized dynamical systems [chen2018neural]. However, in contrast to neural differential equations we learn a symbolic, interpretable representation of the state equation that implements the policy.
@@
In this paper, we introduce a method for evolving dynamic symbolic policies with GP to solve dynamic control tasks (Fig. 1). Using this approach, dynamic symbolic policies are successfully evolved in linear and non-linear environments, under several challenging circumstances such as partial observability and varying environment settings where static policies obtain limited performance. High-performing symbolic policies are still discovered when the control dimension increases, showing that our method is scalable. Furthermore, we find that the evolved symbolic policies obtain comparable performance as the black-box baseline, given by neural differential equations [chen2018neural]. The key contribution of our work is a method for discovering interpretable high-performing dynamic symbolic policies with GP without requiring human expertise.
@@ 2 Methods
The goal of this paper is to evolve dynamic symbolic policies that obtain high performance in control problems. We compare dynamic symbolic policies with static symbolic policies, which are explained in more detail in this section. Afterwards, the specifics of the GP algorithm and how we extended GP to learn multiple trees efficiently are described. In subsequent subsections the baselines and experiments we performed are detailed. The code including our GP implementation, environments and baselines is available at: https://github.com/sdevries0/MultiTreeGP.
Figure 2: Structure and optimisation of the static and dynamic symbolic policies. (a) The feedback loop of an environment coupled with a control system that implements a static policy. C represents the number of control signals required to interact with the environment. For every function $g_{j}$ a tree is optimised. (b) The control system is extended to include a latent state to implement a dynamic policy. H represents the number of units in the latent state. (c) An example representation of how a dynamic symbolic policy is represented by a set of trees. A tree is learned for every latent state and output signal. The functions $f_{i}$ and $g_{j}$ match the control system’s structure defined in panel b. The trees are optimised with genetic programming. (d) In genetic programming, crossover is applied to a pair of parents and produces two offspring individuals. The blue subtrees are interchanged between the two trees. (e) Mutation changes one individual and results in a single offspring. In this example, the blue node is a function node that is changed to a different type of function node.
@@ 2.1 Environment model
To evaluate a candidate symbolic policy proposed by GP, a coupled system consisting of the policy and an environment is simulated, as described in Fig. 2a and Fig. 2b. The environment is modeled by a state $\mathbf{x}\in\mathbb{R}^{N}$, following the dynamics of a stochastic differential equation
@@
The controller receives observations $\mathbf{y}\in\mathbb{R}^{M}$ from the state $\mathbf{x}$, determined by function $g_{\textrm{env}}$ and possibly corrupted by noise. In our experiments, the noisy observations are generated by
@@ 2.2 Control model
Dynamic policies
We extend this approach to dynamic symbolic policies (Fig. 2b), which include memory by using a latent state. The latent state processes incoming observations, and the readout function maps the latent state to a control signal. The dynamic policies contain expressions for the latent dimension and control dimension (Fig. 2c). The latent state is governed by a system of ordinary differential equation, according to
@@
for $1\leq i\leq H$. Note that the dynamic policies also receive the previous control as an additional input. The readout functions of the dynamic policies are similar to Equation 3, but the input to $g_{j}$ is given by the latent state $\mathbf{a}$ instead of the observations $\mathbf{y}$. That is,
@@ 2.4 Genetic programming
The symbolic policies are evolved via genetic programming (GP). GP is a variant of evolutionary algorithms that focuses on learning computer programs [koza1994genetic]. A population of solutions is evolved to satisfy a certain goal or task through stochastic optimisation. In our setup, individuals are represented by tree-structured computational graphs, constructed from predefined sets of function nodes and leaf nodes. Function nodes cover mathematical operators and functions, while leaf nodes describe variables and constants. GP is largely defined by two hyperparameters. The first hyperparameter is the number of generations, which determines how many iterations of evolution are performed. The second hyperparameter is the population size, which defines the number of candidates that are evolved at every generation. See Algorithm 1 for an overview of the GP algorithm. The GP algorithm consists of initialisation, evaluation and reproduction steps, as described in more detail in the following sections.
@@ 2.4.2 Evaluation
To evaluate an individual, its tree is transformed into a program of desired functionality and tested on a problem. The performance of the individual is expressed with a fitness score, computed with a fitness function. The fitness scores are used to select individuals for reproduction, where fitter individuals produce more offspring. After evaluation, parents are selected for reproduction using tournament selection [fang2010review].
@@ 2.4.3 Reproduction
In GP, the new population is built with new candidate solutions that are generated with crossover and mutation. In crossover the genotype of two individuals are combined to produce new trees. Fitter individuals are more likely to be selected for reproduction, therefore the offspring is built from promising tree structures. In both parents a random node from the tree is selected, after which these nodes and their subtrees are swapped. An example of the crossover operator is shown in Fig. 2d. Uniform crossover is a variant of the standard crossover where multiple nodes can be swapped between a pair of trees without swapping full subtrees [spears1991study]. Uniform crossover improves the locality of the crossover operation, but both types of crossover can produce fit individuals and are therefore both included in our implementation.
Mutation was originally not integrated into GP [koza1994genetic], but mutation has shown to be a useful addition to the algorithm [luke1997comparison]. Mutation is applied to a single individual to generate one offspring and is applied instead of crossover in our implementation. Many aspects of trees can be mutated, for example changing, deleting or adding operators, replacing subtrees or changing variables or constants. Fig. 2e shows an example of mutation, in which an operator is changed to a different type of operator. Besides crossover and mutation, the new population also consists of randomly generated individuals and the fittest individuals from the previous generation.
The number of trees that has to be optimised differs for the static and dynamic policies. The static policies evolve a tree for every control input the problem requires. These trees also have to be evolved in the dynamic policies, but additional trees are optimised for every latent state variable. We chose to fix the size of the latent state in the dynamic policies and allow the readout function to learn which latent states to use. To learn both a symbolic latent state and readout function simultaneously in the dynamic policies, the individuals in the GP algorithm are extended to learn multiple trees [cao2000evolutionary]. During reproduction, crossover and mutation are not applied to every tree in the individual to be able to make small jumps through the search space. Crossover is only applied to trees in the same position in two parents, and a third type of crossover is introduced in which only complete trees are swapped between individuals.
@@ 2.4.4 Regularisation
Regularisation is necessary in GP to produce interpretable trees and improve the efficiency of the optimisation. To regularise the size and complexity of the reproduced trees, a punishment relative to the number of nodes in a tree is added to the fitness [luke2006comparison]. Additionally, the depth of the trees is limited to seven, and during initialisation the depth is set to four. To introduce more compact trees in the population, a small part of the population is simplified mathematically at every generation [hooper1996improving]. To improve diversity in the population during evolution, multiple subpopulations are evolved independently [fernandez2003empirical]. By configuring the hyperparameters in each subpopulation differently, exploration and exploitation is enforced explicitly [herrera2000gradual]. Finally, dynamic policies with constant readout functions that only consist of targets or scalars and ignore the latent state, often form local minima because they offer stable solutions. However such policies will never come close to high-performing policies. To improve the efficiency of the evolutionary search, every readout function of a dynamic policy that is reproduced or sampled will have at least one reference to a latent variable to avoid such local minima, or else it is rejected for the next population.

### v5

@@ Abstract
Artificial intelligence techniques are increasingly being applied to solve control problems, but often rely on black-box methods without transparent output generation. To improve the interpretability and transparency in control systems, models can be defined as white-box symbolic policies described by mathematical expressions. For better performance in partially observable and volatile environments, the symbolic policies are extended with memory represented by continuous-time latent variables, governed by differential equations. Genetic programming is used for optimisation, resulting in interpretable policies consisting of symbolic expressions. Our results show that symbolic policies with memory compare with black-box policies on a variety of control tasks. Furthermore, the benefit of the memory in symbolic policies is demonstrated on experiments where memory-less policies fall short. Overall, we present a method for evolving high-performing symbolic policies that offer interpretability and transparency, which lacks in black-box models.
@@ 1 Introduction
Many problems in our increasingly complex society can be viewed as control problems, where the goal is to make optimal decisions in real time such as to regulate the behaviour of the controlled system, ranging from smart infrastructure systems to healthcare systems [Annaswamy2023]. In recent years, there have been many successful applications of artificial intelligence (AI) to control problems, offering automated solutions to regulate complex non-linear processes. While the employed control algorithms achieve high performance, they usually require the use of black-box models such as neural networks [adadi2018peeking]. Black-box models often allow for highly general computations, but lack interpretability of their functioning, making it difficult to understand the computed decisions and actions [loyola2019black, glanois2024survey]. Transparency is particularly important in control systems, because it aids in identification of biases, fault detection and improvement and could therefore raise confidence of users in the system’s reliability [zhang2021survey].
@@
In contrast to black-box models, white-box models allow for easy inspection of their internal mechanisms [adadi2018peeking]. However, white-box models show an accuracy-interpretability trade-off and there are currently no general training methods to produce white-box approaches that can compete with black-box approaches in terms of efficiency and accuracy [loyola2019black, puiutta2020explainable]. Most white-box approaches opt for post-training model distillation to produce a transparent model that learns the input-output mapping of a black-box model [gou2021knowledge]. Disadvantages of model distillation are that it requires a two-phase learning process and information loss could occur in the distillation phase [ras2022explainable]. Therefore, it may be more desirable to learn white-box models directly. When optimising an inherently white-box model, interpretability of decisions already occurs at intermediate stages, showing transparent learning, and outputs are directly dependent on explainable features [ras2022explainable].
@@
A promising candidate for white-box control systems are symbolic policies [calegari2020integration, landajuela2021discovering, gou2021knowledge]. Symbolic policies are described by interpretable symbolic expressions, consisting of mathematical operators and state variables. Symbolic policies do not require a predefined model structure, therefore it is less affected by human bias or knowledge gaps compared to other data-driven approaches. Additionally, symbolic policies may capture rules that generalise better to other settings, and learned equations tend to be more efficient than neural networks in terms of parameter count [landajuela2021discovering].
@@
Due to the non-differentiable structure of mathematical expressions, gradient-based methods cannot easily be applied to the task of the optimising symbolic policies. A viable alternative is genetic programming (GP), an evolutionary algorithm that automatically learns computer programs [koza1994genetic]. GP has seen great successes in symbolic regression problems, which seek to discover interpretable mathematical models describing data [cao2000evolutionary, sakamoto2001inferring, iba2008inference, schmidt2009distilling, quade2016prediction, bongard2007automated, wang2019symbolic]. Therefore we consider GP as a promising approach to optimise symbolic policies to solve control tasks [hein2018interpretable, nadizar2024naturally].
Current approaches to symbolic policy learning focus primarily on memory-less policies that have a fixed mapping from observations to control outputs [guo2024efficient, hein2018interpretable, nadizar2024naturally, landajuela2021discovering, gou2021knowledge, shimooka2000generating]. However, such policies may apply poorly to real-world settings, where volatility or partial observability complicate tasks. By adding memory to a control policy, past observations to improve policies in difficult settings [heess2015memory, khan2017memory]. This improvement has also been found by extending symbolic policies with memory [koza1992evolution, koza1999genetic, andersson1999reactive, raik1997evolving], where GP explicitly learns when to store and retrieve observations from memory.
Alternatively, memory can be defined by dynamic variables that continuously integrate observations [rackauckas2020universal]. By modelling these variables with differential equations, the policy can use them to estimate unobserved variables or learn other important properties of the environment. This is similar to the application of neural differential equations to estimate parameterised dynamical systems [chen2018neural]. However, in contrast to neural differential equations, the latent variables in a symbolic policy are represented by symbolic equations, which makes the policy more transparent. Besides the latent variables, the policy also consists of a symbolic readout function that computes a control signal given the latent variables.
@@
In this paper, we introduce a method for evolving memory-based symbolic policies with GP to solve dynamic control tasks, where continuous-time dynamically changing latent variables represent the memory (Fig. 1). Using this approach, symbolic policies are successfully evolved in linear and non-linear environments, under several challenging circumstances such as partial observability and varying environment settings, where memory-less policies obtain limited performance. High-performing symbolic policies are still discovered when the control dimension increases, showing that our method is scalable. Furthermore, we find that the evolved symbolic policies obtain performance comparable to the black-box baseline, given by neural differential equations [chen2018neural]. The key contribution of our work is a method for discovering interpretable high-performing memory-based symbolic policies using GP without requiring human expertise.
2 Related Work
Already in the early days of GP, it was proposed to find policies to solve control tasks [koza1994genetic, koza2000automatic]. These symbolic policies map the incoming observations to a control signal that was fed into the environment. Hereafter, GP has seen several applications for learning symbolic control policies [shimooka2000generating, hein2018interpretable, nadizar2024naturally], although differentiable methods have also proven to be successful at finding symbolic policies [landajuela2021discovering, gou2021knowledge]. Furthermore, GP has been applied to the optimisation of decision trees [custode2023evolutionary] or fuzzy controllers [hein2018generating] to solve control tasks. Other evolutionary algorithms have shown to optimise control policies that compete with gradient-based approaches [salimans2017evolution, stanley2002evolving, such2017deep].
Before applying control policies to more realistic settings, the policies have to be able to handle partial observability and volatile environments. In [hein2018interpretable], the policy receives observations from past time points, which enables the policy to control a stochastic and partially observable task, although the policy itself does not contain explicit memory. This method requires storing a large number of observations, which does not scale well to longer trajectories or higher dimensions [khan2017memory], as the search space of GP becomes very large. A more general approach is to allow GP to learn to store computed values, which may be used for future computations [koza1999genetic, koza1992evolution, raik1997evolving, andersson1999reactive]. In [kelly2023discovering], control algorithms with explicit memory storage are evolved with genetic algorithms, and the resulting policies are able adapt to changes in the environment.
In contrast to these approaches, in this paper we will use GP to evolve the state equations of continuous-time latent variables that function as memory. Incoming observations are integrated into the dynamics of the latent variables, which offers a more general and scalable approach than using time-lagged observations [khan2017memory]. Instead of explicitly writing or reading values to memory, the latent variables are continuously updated by new observations and are added as an input to a readout that computes the control signal. This way, the policy can automatically select the relevant features to efficiently estimate hidden states, whereas using time-lagged observations would increase the number of irrelevant variables. The latent memory also improves estimation of variables from stochastic observations or adaptation environment settings. For clarity, in this paper the continuous-time memory-based policies are referred to as dynamic policies, and the memory-less policies as static policies.
The dynamic policies consist of multiple latent variables and a readout function, therefore multiple trees have to be optimised with GP simultaneously [langdon1998genetic]. In [cao2000evolutionary], GP evolves multiple trees that together represent a system of coupled differential equations. Although the trees could be learned independently [bongard2007automated], this setup is required when certain variables are not observed [cornforth2013inference]. This representation of multiple trees has also been utilised in collective decision making making [wen2016learning, brameier2001evolving] and multi-class classifiers [muni2004novel, ingalalli2014multi]. Based on these works, we extend our GP implementation to include tree based crossover and only apply crossover to pairs of trees from the same class.
@@ 3 Methods
The goal of this paper is to evolve dynamic symbolic policies that obtain high performance in control problems. We compare dynamic memory-based symbolic policies with static memory-less symbolic policies, which are explained in more detail. Afterwards, the specifics of the GP algorithm and how we extended GP to learn multiple trees efficiently are described. In the subsequent subsections the baselines and experiments we performed are detailed. For the implementation of GP, we utilised Kozax [^1] https://github.com/sdevries0/Kozax, a recently developed GP framework written in JAX. The code for reproducing the experiments and baselines is available at: https://github.com/sdevries0/DSP_paper.
Figure 2: Structure and optimisation of the static and dynamic symbolic policies. (a) The feedback loop of an environment coupled with a control system that implements a policy without memory. C represents the number of control signals required to interact with the environment. For every function $g_{j}$ a tree is optimised. (b) The control system is extended to include latent variables to add memory to the policy. H represents the number of latent variables. (c) An example representation of how a symbolic policy with latent variables is represented by a set of trees. A tree is learned for the state equation of every latent variable and output signal. The functions $f_{i}$ and $g_{j}$ match the control system’s structure defined in panel b. The trees are optimised with genetic programming. (d) In genetic programming, crossover is applied to a pair of parents and produces two offspring individuals. The blue subtrees are interchanged between the two trees. (e) Mutation changes one individual and results in a single offspring. In this example, the blue node is a function node that is changed to a different type of function node.
@@ 3.1 Environment model
To evaluate a policy, a coupled system consisting of the policy and an environment is simulated, as described in Fig. 2a and Fig. 2b. In our paper, the policies are evaluated on dynamic tasks modelled as Markov decision processes. The environment is modelled by a state $\mathbf{x}\in\mathbb{R}^{N}$, following the dynamics of a stochastic differential equation
@@
The policy receives observations $\mathbf{y}\in\mathbb{R}^{M}$ from the state $\mathbf{x}$, determined by function $g_{\textrm{env}}$ and possibly corrupted by noise. In our experiments, the noisy observations are generated by
@@ 3.2.1 Static policies
3.2.2 Dynamic policies
We extend this approach to dynamic symbolic policies (Fig. 2b), which implement memory through continuous-time dynamic latent variables. The latent variables process incoming observations, and the readout function maps the latent variables to a control signal. The dynamic policies contain expressions for the latent dimension and control dimension (Fig. 2c). The latent variables are governed by a system of ordinary differential equation, according to
@@ 3.2.2 Dynamic policies
for $1\leq i\leq H$. Note that the dynamic policies also receive the previous control as an additional input. The readout functions of the dynamic policies are similar to Equation 3, but the input to $g_{j}$ is given by the latent variables $\mathbf{a}$ instead of the observations $\mathbf{y}$, which results in
@@ 3.4 Genetic programming
The symbolic policies are evolved via genetic programming. GP is a variant of evolutionary algorithms that focuses on learning computer programs [koza1994genetic]. A population of solutions is evolved to satisfy a certain goal or task through stochastic optimisation. In our setup, individuals are represented as tree-structured computational graphs, constructed from a predefined set of function nodes and leaf nodes. Function nodes cover mathematical operators and functions, while leaf nodes describe variables and constants. GP is largely defined by two hyperparameters. The first hyperparameter is the number of generations, which determines how many iterations of evolution are performed. The second hyperparameter is the population size, which defines the number of candidates that are evolved at every generation. See Algorithm A1 for an overview of the GP algorithm. The GP algorithm consists of initialisation, evaluation and reproduction steps, as described in more detail in the following sections.
@@ 3.4.2 Evaluation
To evaluate a candidate policy, the trees are transformed into callable programs and tested on a problem. The performance of the individual is expressed with a fitness score, computed with a fitness function. The fitness scores are used to select individuals for reproduction, where fitter individuals produce more offspring. After evaluation, parents are selected for reproduction using tournament selection [fang2010review].
@@ 3.4.3 Reproduction
Every generation a new population is generated, consisting of candidate solutions that are generated with crossover and mutation. In crossover the genotype of two individuals is combined to produce new trees. Fitter individuals are more likely to be selected for reproduction, therefore the offspring is built with promising tree structures. In both parents a random node from the tree is selected, after which these nodes and their subtrees are swapped. An example of the crossover operator is shown in Fig. 2d. Uniform crossover is a variant of the standard crossover where multiple nodes can be swapped between a pair of trees without swapping full subtrees [spears1991study]. Uniform crossover improves the locality of the crossover operation, but both types of crossover can produce fit individuals and are therefore both included in our implementation.
Mutation was originally not integrated into GP [koza1994genetic], but mutation has shown to be a useful addition to the algorithm [luke1997comparison]. Mutation is applied to a single individual to generate one offspring and is applied instead of crossover in our implementation. Many aspects of trees can be mutated, for example changing, deleting or adding operators, replacing subtrees or changing variables or constants. Fig. 2e shows an example of mutation, in which an operator is changed to a different type of operator. Besides crossover and mutation, the new population also consists of randomly generated individuals and the elite individuals from the previous generation.
The number of trees that has to be optimised differs for the static and dynamic policies. The static policies evolve a tree for every control input the problem requires. These trees also have to be evolved in the dynamic policies, but additional trees are optimised for every latent variable. We chose to fix the dimensionality of the latent state in the dynamic policies and allow the readout function to learn which latent variables to use. To learn both a symbolic latent state and readout function simultaneously in the dynamic policies, the individuals in the GP algorithm are extended to learn multiple trees [cao2000evolutionary]. During reproduction, crossover and mutation are not applied to every tree in the individual to be able to make small jumps through the search space. Crossover is only applied to trees in the same position in two parents, and a third type of crossover is introduced in which only complete trees are swapped between individuals [muni2004novel].
@@ 3.4.4 Regularisation
Regularisation is necessary in GP to produce interpretable trees and improve the efficiency of the optimisation. To regularise the complexity of the reproduced trees, a punishment relative to the number of nodes in a tree is added to the fitness [luke2006comparison]. Additionally, the depth of the trees is limited to seven, and during initialisation the depth is set to four. To introduce more compact trees in the population, a small part of the population is simplified mathematically at every generation [hooper1996improving]. To improve diversity in the population during evolution, multiple subpopulations are evolved independently [fernandez2003empirical]. By configuring the hyperparameters in each subpopulation differently, exploration and exploitation is enforced explicitly [herrera2000gradual].